OpenStudy (rosedewittbukater):

Can someone help me find the domain and range of y=3csc(3x+pi)-2 ?

OpenStudy (anonymous):

idk because Im in 7th grade...

OpenStudy (rosedewittbukater):

@ganeshie8 Can you help me with this math problem?

OpenStudy (rosedewittbukater):

@e.mccormick @thomaster Can you help me with this math problem? (AP Calculus, Trig review)

OpenStudy (rosedewittbukater):

I know that the domain is supposed to be kpi/3, but I don't know how to get that.

OpenStudy (rosedewittbukater):

@e.mccormick

OpenStudy (rosedewittbukater):

@thomaster

OpenStudy (rosedewittbukater):

@amistre64 @ganeshie8 Can you help me?

OpenStudy (amistre64):

the csc function rides on the humps of which function?

OpenStudy (amistre64):

whenever the function is rides on goes zero, the inverse hits a divide by zero and we eliminate those from the domain

OpenStudy (amistre64):

the 'parent' functions range llives between a min and max, the inverse hits everything else everything else

OpenStudy (amistre64):

in other words, exploring the parent function will help us determine this one

OpenStudy (rosedewittbukater):

@amistre64 so how do I do that to find the domain?

OpenStudy (amistre64):

what is the 'parent' function that csc is an inverse of?

OpenStudy (rosedewittbukater):

y=sin x ?

OpenStudy (amistre64):

inverse is a bad name for it in trig .... reciprocal and yes, sin of whtever argument they used

OpenStudy (amistre64):

now, there is only one real number that doesnt have a good reciprocal: 0 if we flip it we get n/0 soo ... we need to exclude every argument that makes sin go zero

OpenStudy (rosedewittbukater):

How do I find what makes sin zero?

OpenStudy (amistre64):

um, that is going to be one of those basic trig questions that you should have in memory already

OpenStudy (amistre64):

sin(u) = 0 then solve the argument they give you to equate it to u

OpenStudy (rosedewittbukater):

What is u?

OpenStudy (amistre64):

u is an argument that we can use that is not x ....

OpenStudy (rosedewittbukater):

I feel like I should know all of this, but I don't remember learning it. The textbook doesn't say anything about it either.

OpenStudy (amistre64):

when does sin equal zero? you need to know this, ill correct you if need be but i cant just tell you outright

OpenStudy (rosedewittbukater):

You mean on the unit circle?

OpenStudy (amistre64):

i cant draw at the moment or id demonstrate a unit circle

OpenStudy (amistre64):

yeah

OpenStudy (rosedewittbukater):

Everyone else online was saying you have to solve for sin (3x+ pi) = k pi or something like that so maybe that's confusing me but I'm not sure.

OpenStudy (amistre64):

itll be more like, find when sin(u) = 0 then u = 3x+pi will solve for x (u - pi)/3 = x

OpenStudy (rosedewittbukater):

I'm pretty sure its zero at 0 and pi

OpenStudy (amistre64):

0 and pi are good, but aslo 2pi, 3pi, 4pi .... every integer multiple of pi makes sin go zero

OpenStudy (amistre64):

so, lets say u = k pi, such that k is an integer

OpenStudy (rosedewittbukater):

okay

OpenStudy (amistre64):

so, solving for x, as i already did (k pi-pi)/3 = x has to be excluded from the set of Real numbers and that defines our domain

OpenStudy (amistre64):

(k-1)pi/3 but since k is an integer, then k-1 is also any integer lets say every k pi/3 for brevity

OpenStudy (amistre64):

\[D=\{x \in R:x\ne k\frac{\pi}{3}, k\in Z\}\]

OpenStudy (amistre64):

now the range is simple enough, at least for me since i already know my trig :) its the center line, plus/minus the amplitude

OpenStudy (rosedewittbukater):

(-infinity, -5) U (1, infinity) ?

OpenStudy (amistre64):

include the -5 and 1

OpenStudy (amistre64):

sin=csc at the humps

OpenStudy (rosedewittbukater):

What do you mean by include

OpenStudy (amistre64):

your notation, when we want to omit an endpoint we close it with ) when we want to include it, it EQUALS it, we draw an equal sign connected by a bar: ]

OpenStudy (amistre64):

since infinity can never be equated, you use ( and ) which is fine, but -5 and 1 are PART of the range and need to be included since we can equal -5 or 1

OpenStudy (rosedewittbukater):

Oh okay so just use brackets instead of parentheses?

OpenStudy (amistre64):

for the -5 and 1 yes

OpenStudy (rosedewittbukater):

ok thank you. There's another problem like this but it uses tangent instead of csc. Do you think you could help me out on that as well?

OpenStudy (amistre64):

tangent is a common trig function, its built out of sin and cos: tan = sin/cos so when do you think we would omit a part of the domain?

OpenStudy (rosedewittbukater):

At pi/2 and 3pi/2?

OpenStudy (rosedewittbukater):

That's when cot is zero I believe

OpenStudy (amistre64):

odd integer multiples of pi/2 yep

OpenStudy (rosedewittbukater):

the equation is y = -3 tan (2x + pi) + 2

OpenStudy (rosedewittbukater):

Not sure what to do next..

OpenStudy (rosedewittbukater):

oops I meant y = -3 tan (3x + pi) + 2

OpenStudy (rosedewittbukater):

3x + pi = kpi/2 ?

OpenStudy (rosedewittbukater):

I'll take that as a yes..